The following information is taken from the 2020 annual report of Wildhorse, Inc. Wildhorse’s fiscal year ends December 31 of each year. Wildhorse’s December 31, 2020, balance sheet is as follows.
Wildhorse, Inc.
Balance Sheet
December 31, 2020
Assets
Cash $420
Inventory 1,770
     Total current assets 2,190
Plant and equipment 1,800
     Accumulated depreciation (144)
       Total assets 3,846
Liabilities
Bonds payable (net of discount) $1,426
Stockholders’ equity
Common stock 1,500
Retained earnings 920
       Total liabilities and stockholders’ equity $3,846

Note X: Long Term Debt:
On January 1, 2019, Wildhorse issued bonds with face value of $1,500 and a coupon rate equal to 10%. The bonds were issued to yield 12% and mature on January 1, 2024.

Additional information concerning 2021 is as follows.
1. Sales were $3,400, all for cash.
2. Purchases were $2,000, all paid in cash.
3. Salaries were $640, all paid in cash.
4. Property, plant, and equipment was originally purchased for $1,800 and is depreciated straight-line over a 25-year life with no salvage value.
5. Ending inventory was $1,900.
6. Cash dividends of $110 were declared and paid by Wildhorse.
7. Ignore taxes.
8. The market rate of interest on bonds of similar risk was 12% during all of 2021.
9. Interest on the bonds is paid semiannually each June 30 and December 31.

Prepare an income statement for the year ending December 31, 2021. Assume semiannual compounding of the bond interest.

Prepare a balance sheet for Wildhorse, Inc. at December 31, 2021. Assume semiannual compounding of the bond interest.

Compute debt to assets ratio and times interest earned ratio for 2020 and 2021. Wildhorse’s net income in 2020 was $590 and interest expense was $169.
Answer #1

3,400 1,870 1,530 Wildhorse, Inc. Income Statement For the Year Ending December 31, 2021 Sales Cost of goods sold Gross profi

Wildhorse, Inc. Balance Sheet As of December 31, 2021 Assets Current Assets Cash Inventory Total Current Assets 920 1,900 2,8

Wildhorse, Inc. 2020 37.08% 4.49 Calculation 2021 2020 2021 Formula 32.87% =1,426/3,846 =1,448/4,404 Total Debt / Total Asset

+ Wildhorse, Inc. Assets Liabilties + Equity Account Titles for Events Cash Inventory Plant & Equip Accumulated Dep = Bonds PWildhorse, Inc. Schedule of Bond discount Amortization Calculation Carrying Amount Interest Payment Date Interest Payment Int